OpenStudy (dls):

Inverse trigonometry question doubt. \LARGE f(x)=\cos^{-1}x+\cos^{-1}(\frac{x}{2}+\1/2 sqrt{3-3x^2})

OpenStudy (dls):

\[\LARGE f(x)=\cos^{-1}x+\cos^{-1}(\frac{x}{2}+\sqrt{3-3x^2})\] Then A)f(2/3)=pi/3 B)f(2/3)= 2 arccos(2/3)-pi/3 C)f(1/3)=pi/2 D)f(1/3)=2 arccos(1/3)-pi/3

OpenStudy (dls):

I tried to solve the 2nd part by substituting x=sin theta It gave me, \[\cos^{-1}(\frac{1}{2}\sin \theta+\frac{\sqrt{3}}{2}\cos \theta)=\cos^{-1}(\sin(\frac{\pi}{3}+\theta))=\frac{\pi}{2}-\sin^{-1}(\sin(\frac{\pi}{3}+\theta)\]

OpenStudy (dls):

sorry thats should be (pi/6+theta)

OpenStudy (dls):

\[=>\frac{\pi}{2}-\frac{\pi}{6}-\theta=\frac{\pi}{3}-\sin^{-1}x\]

OpenStudy (anonymous):

why do you think it's \(\pi/6\)?

OpenStudy (dls):

hmm its not :P

OpenStudy (dls):

its sin ( pi/3+theta) only

OpenStudy (dls):

i got confused with smt else ,nvm.

OpenStudy (dls):

\[\LARGE \frac{\pi}{2}-\frac{\pi}{3}-\theta=\frac{\pi}{6}-\sin^{-1}x\]

OpenStudy (dls):

\[\Large =>\cos^{-1}x-\sin^{-1}+\frac{\pi}{6}=\cos^{-1}-(\frac{\pi}{2}-\cos^{-1}x+\frac{\pi}{6})\]

OpenStudy (dls):

cos^-1x*

OpenStudy (dls):

\[\LARGE 2\cos^{-1}x- \frac{2 \pi}{3}\] is this much correct?

OpenStudy (dls):

@oldrin.bataku

OpenStudy (dls):

can anybody hear me...or am I talking to myself..

OpenStudy (dls):

my mind is running empty..in this search for someone else..

OpenStudy (dls):

who doesn't look right through me..its all just static in my head..

OpenStudy (dls):

can anybody tell me..why..im lonely like a satellite..

OpenStudy (anonymous):

did you mean$$\LARGE f(x)=\cos^{-1}x+\cos^{-1}(\frac{x}{2}+\frac12\sqrt{3-3x^2})$$

OpenStudy (dls):

that is what i wrote?

OpenStudy (anonymous):

no this has a \(1/2\) before the radical

OpenStudy (dls):

oh sorry 1/2 is missing sorry,1/2 is there too

OpenStudy (anonymous):

Substituting \(x=\cos t\):$$\begin{align*}f(\cos t)&=t+\arccos\left(\frac12\cos t+\frac{\sqrt3}2\sin t\right)\\&=t+\arccos(\sin(\pi/6+t))\\&=t+\arccos(\cos(\pi/2+\pi/6+t))\\&=2t+2\pi/3\end{align*}$$for \(t\in[0,\pi/3)\) hence for \(x\in(1/2,1]\) we have:$$f(x)=2\arccos(x)+2\pi/3\\f(2/3)=2\arccos(2/3)+2\pi/3$$ which is neither option

OpenStudy (anonymous):

ok so now we continue

OpenStudy (dls):

so i did correct?

OpenStudy (dls):

oh a difference of sign :O

OpenStudy (ybarrap):

using f(2/3) I get one of your options.